KyroKazimake:

I'm in math class and I'm sort of confused. Isn't 1-cosx= Sinx

KyroKazimake:

@Hero @Shadow

ThisGirlPretty:

@Nnesha

Shadow:

\[\sin^2 x + \cos ^2 x = 1\]

Shadow:

Talking about this identity?

KyroKazimake:

@Shadow yes

Shadow:

\[\sin \theta = \sqrt (1 - \cos \theta)\]

KyroKazimake:

Ohhhhhh ok I see now.

KyroKazimake:

Thanks shadow